OpenStudy (anonymous):

Solve the system, using substitution. What is the n coordinate of the solution? 2m - 3n = 6 and m = 3 - 2n A. 2 B. 0 C. -4 D. 2 E. no solution

OpenStudy (luigi0210):

Plug them in and solve... 2(3-2n)-3n=6 6-4n-3n=6 -7n=0 n=0

OpenStudy (anonymous):

2m - 3n = 6 and m = 3 - 2n 2m - 3n = 6 2(3 - 2n) - 3n = 6 6 - 4n - 3n = 6 . . . .

OpenStudy (mayankdevnani):

@micahcarr76 any confusion??